OpenStudy (anonymous):

Find the thirtieth term of the following sequence. -6, -4, -2, 0, ...

OpenStudy (anonymous):

if you want to do it the easy way... you're adding 2 every term to what is -6 + 12*(2)?

OpenStudy (anonymous):

you multiply by 12 because you start with the first term, and you need to find the value 12 terms later for the 13th term

OpenStudy (kropot72):

The formula needed for this arithmetic progression is \[n ^{th}\ term=a+(n-1)d\] So the thirtieth term is \[-6+(30-1)2=(2 \times 29)-6\]

OpenStudy (anonymous):

i just want the answer im in a exam

OpenStudy (anonymous):

it only give this options 52 64 66

OpenStudy (kropot72):

Just calculate the following: \[(2 \times 29)-6=you\ can\ calculate\]